# CoPlay: Audio-agnostic Cognitive Scaling for Acoustic Sensing

### v1

@@ 3.1. Understanding the effect of dynamically scaling the signal
Achieving our two objectives simultaneously is hard because just tweaking the signal amplitude in the time domain will cause noisy distortion in the frequency domain. To understand the effect of inordinately tweaking the amplitude, consider amplitude modulation (AM) as an interpretable example. AM is a technique used in telecommunications to transmit information by varying the amplitude of a carrier wave. In Fig. 2, the high-frequency carrier is multiplied by the low-frequency base signal, i.e. amplitude modulated. It causes the frequency spectrum of the signal to spread out with noise. This is because the noisy side bands, located above and below the carrier frequency, are created by the modulation process. In this case, the change is closed-form; the side frequencies are $f_{c}+f_{m}$ and $f_{c}-f_{m}$. (Derivatives can be found in Appendix A). Building upon this observation, our cognitive scaling is namely multiplying the signal with a more complex window function, leading to various sidebands in the frequency. Therefore, this effect of altering signals necessitates the utilization of deep learning in this work; closed-form signal processing methods or traditional machine learning can hardly dynamically control the distortion during generation.
Figure 2. Inordinate scaling in time domain causes distortion in frequency domain, taking Amplitude Modulation as an example.
3.2. Sensing signal modulation types
In the field of acoustic sensing, two commonly used types of signals are sine waves and FMCW chirps. These signals serve different purposes and are utilized in various applications within acoustic sensing systems.
@@ 3.2. Sensing signal modulation types
Sine waves: A sine wave is a single-frequency oscillation that varies with time according to the mathematical function $A\sin(2\pi ft+\phi)$, where $A$ is the amplitude of the sine wave, $f$ is the frequency, $\phi$ is the initial phase. They are commonly used in acoustic sensing for tasks such as signal testing, calibration, and as reference signals. They provide a simple and well-defined waveform that is easy to generate and analyze.
FMCW chirps: Frequency Modulated Continuous Wave (FMCW) is a technique where the frequency of a continuous wave changes linearly over time. Basically, they are repeated chirps defined as $A\sin(2\pi(f_{a}t+\frac{f_{1}-f_{0}}{2T}t^{2}))$, where $f_{0}$ and $f_{1}$ are the starting and ending frequency of the chip; $T$ is the time duration of one chirp. They are widely used in applications like radar and sonar. In acoustic sensing, FMCW signals offer advantages such as improved range resolution and the ability to distinguish between multiple targets. They are particularly useful in applications requiring precise distance measurements and target discrimination.
3.3. Define the problem as optimization
@@
Input and output: Let $x$ be the sensing signal, i.e. a high-frequency sine or chirp. Concurrently, an App plays music $z$. Then the mixer mixes them within the maximum range of $[-m,m]$ allowed by the number of DAC bits. In this paper, we use signed 16-bit PCM, resulting in a range of $-2^{15}\sim 2^{15}-1$. For computation convenience, we set $m=2^{15}-1$. In general, our goal is to generate a cognitively-scaled sensing signal $\hat{x}$ conditioned on the music $z$; and the $\hat{x}$ is optimal when satisfying the following objectives defined in the loss.
Loss: The first objective is to maximize the sensing amplitude within the bandwidth left by music. In other words, the amplitude of the mixed signal $\hat{x}+z$ in time domain should be close to the maximum allowed amplitude $m$. To restate, it is to minimize loss $q(\hat{x}+z,m)$, namely the difference between the magnitude of them. However, this would cause consequential frequency distortion as explained in section 3.1, which deteriorates the receiver-side algorithms since they generally rely on accurate frequency (either FFT convolution or correlation convolution). Therefore, we define another loss $p(x,\hat{x})$ as the difference between $x$ and $\hat{x}$ in frequency domain; it is the difference between their L-2-norm of FFT coefficients. In summary, our goal is to minimize these two loss terms simultaneously as an optimization problem.
@@
(1) $p(x,\hat{x})=1-\|\hat{c_{f_{c}}}\|_{2}+\frac{1}{N-1}\sum_{i\in-{f_{c}}}\|c_{i}-\hat{c_{i}}\|_{2}$
where $c$ are the magnitude of the complex number coefficients in FFT, $FFT(x)=(c_{i},...)$, $FFT(\hat{x})=(\hat{c_{i}},...)$. And $f_{c}$ is the target frequency bin containing the ultrasound frequency. $N$ is half of the window size since we only take the one symmetric half of the FFT output. The magnitude is normalized by the window size and the maximum allowable amplitude of input $x$, i.e. $m$. The normalization guarantees that the two variable terms in $p$ are standardized and unbiased. We call the first half target-loss and the second half recovery-loss. The intention here is to encourage a large value in the target frequency bins, and a close-to-zero value in the others.
Secondly, $q$ is the L-2 norm of $\hat{x}+z$ in time domain normalized by $m$, named as amplitude-loss.
(2) $q(\hat{x}+z,m)=1-\frac{1}{2N}\|\frac{\hat{x}+z}{m}\|_{2}$
For FMCW chirp, we also want the multiple frequency bins to have even amplitude so that it is well correlated. So, we have an additional variance-loss for FMCW using the standard deviation with zero correction:
(3) $s(\hat{x})=\sqrt{\frac{\sum_{i\in{f_{c}}}(\hat{c_{i}}-\bar{\hat{c_{i}}})^{2}}{N_{f_{c}}-1}}$
@@
(4) $loss_{sine}$ $=\alpha*p(x,\hat{x})+\beta*q(\hat{x}+z,m)$
(5) $loss_{chirp}$ $=\alpha*p(x,\hat{x})+\beta*q(\hat{x}+z,m)+\gamma*s(\hat{x})$
Note that, our loss may not guarantee that $\hat{x}+z$ falls within [$-m,m$]. Therefore, we need to embed this constraint into model layers, as introduced in the following section 3.4 per customized link function.
In practice, we first split the sensing signal and music into small windows and run the optimization for each window. For chirps, the window size should be the size of one chirp. In this paper, we use 18kHz sine wave and 18k-20kHz FMCW signal, with a window size of 512 and a sampling rate of 48kHz. These are common settings and factor in all considerations per efficiency and performance. And our model can be extended to other settings as long as training settings are the same as those in testing.
3.4. Cognitive scaling neural network
@@
Figure 3. Cognitive scaling neural network.
3.4.1. Model architecture with WaveNet backbone
@@
To elaborate, it combines a non-auto-regressive sequence-to-sequence structure with WaveNet backbones. While WaveNet architecture was originally developed for AR modeling, it has since become a basic backbone of various audio models. And we validate this choice in the micro-benchmark section by experimental comparison with sample CNN and RNN backbones. In detail, it employs a stack of dilated convolutional layers which provides better reception fields than conventional ones, making it adept at raw waveform modeling and processing sequential data more efficiently than RNNs. The input of the model is concatenated sensing signal $x$ and music $z$ of the same length of window size 512. Then the 2-row tensor is passed through 2 layers of Wavenet with fused addition and $tanh$ + multiplication, then dilated convolution (hidden size of 2, kernel size of 5, and dilation rate of 1), and finally residual skip layer (double hidden size except the last layer). Then the outputs pass through conv1d layer to merge the two rows, followed by custom layers to be described in the following sub-sections. Based on observation, we choose a large learning rate of 0.1 with Adam optimizer that is stable in training. For implementation, we refer to the Pytorch code base from DiffSinger (Liu et al., 2022) on HuggingFace.
3.4.2. A sinc kernel for anti-aliasing
At the early stage of experiments, we observe non-negligible low-frequency noise in the model output, especially when modeling FMCW chirps. It is a common problem discussed in the computer vision community that the neural network learns low frequency faster and also down-scaling in layers like convolution and max pooling could induce aliasing.
Therefore, we design a custom sinc kernel for band-passing middle activations to isolate specific frequency bands of interest within the signal. The unique properties of the (sinc) filter, notably its brick-wall characteristic in the frequency domain as its Fourier transform is a rectangular function (Appendix  B), makes it a preferred choice of filters over Hamming, etc. So, by putting (sinc) kernel as a layer before the final link function, it mitigates artifacts that may arise in the first stage. For implementation, we use sinc_impulse_response function in the Pytorch experimental nightly version. The negative of it is a high-pass filter. And they can sum to one band-pass filter or more. To demonstrate its effectiveness, we perform ablation in the micro-benchmark section.
Another advantage is its potential to serve as a parameterized kernel when building models for new types of sensing signals other than chirp or sine. A parameterized cutoff could learn a dynamic filter for even multiple signals trained in one model. While we currently train standalone models dedicated to each sensing signal, so that each model achieves optimal performance since we focus on substantiating and validating the underlying concept.
3.4.3. A link function for normalization with agnostic-music as bias
As described, $\hat{x}+z$ should be within the range of [$-m,m$], namely a normalization with agnostic music as bias. To restate, the output $\hat{x}$ should be $a\times m+z$ where $a$ is the activation from the previous layer within [0, 1]. So, to embed this constraint as a trainable layer, we leverage the concept of link function in Generalized Linear Models (GLM); It is in analogy to the logit function for the logistic regression, which uses the prior to target on the probability of binary classification ranging from 0 to 1. Similarly, CoPlay uses a hyperbolic tangent function ($tanh$) as link function that scales the output within [-1, 1]. A bonus is that it encourages the value to approach -1 or 1, which is useful in our cases because we want to maximize the magnitude of the waveform, i.e. close to 1 or -1, as illustrated in Fig. 4. In contrast, clipping simply truncates the output, resulting in loss of information or under-fitting, while $tanh$ is more effective especially when data is continuous and has a wide range of values.
So, we design the $tanh$ layer as follows rather than a min-max clipping: $\hat{x}=tanh(a)*m-z=\frac{e^{a}-e^{-a}}{e^{a}+e^{-a}}*m-z$ where $a$ represents the activations from the previous layer. This scales $\hat{x}+z$ within $[-m,m]$. Moreover, during inference, we feed the $\hat{x}+z$ into the mixer in case $\hat{x}$ exceeds the numerical range. Finally, to quantify the output as integers while the model is trained with float64, our code employs a workaround using the round function: it detaches the difference and then adds it back. Otherwise, the round function or typecast function is not differentiable in PyTorch.
Figure 4. We design a link function using $tanh$ to match the domain and encourage large amplitude. In contrast, $clip$ is only capable of the first.
4. Micro-benchmark
In this section, we detail the experiment setup and assess the model performance under various micro-benchmarks, including evaluating the generation across different datasets and music types, as well as at varying levels of music volume. Additionally, ablation studies are conducted to understand the impact of model architecture and multi-task learning weights. Finally, we offer an interpretable analysis of the output’s quality by visualization.
Evaluation metric: In this section, we mainly use loss as the evaluation metric. The detailed definition of loss is in section 3.3. It is the indicator of the relative performance when conducting micro-benchmark experiments. Furthermore, to better interpret the performance, we visualize the output via their frequency spectrums and cross-correlation ranging profiles. Besides, we will further assess the performance in downstream field studies of respiration and gesture detection in the following section. The code will go public on Github upon the acceptance of the paper.
4.1. Datasets
We use four types of concurrent audio to study model’s generalization across agnostic types of music. They are piano music, hot songs, bass music, and speech datasets. Furthermore, for each type, we select two different datasets.
Piano: One dataset is the Beethoven dataset (Mehri et al., 2016), consisting of piano sonatas. It has a total duration of 10 hours and serves as a benchmark music dataset for audio generation. We interpolate the data to a sample rate of 48kHz and cast the 8-bit quantization into a 16-bit signed integer. Another piano dataset is the YouTubeMix dataset (Samplernn., 2017) with higher-quality recordings, with a total duration of 4 hours. And we format it in the same way as the Beethoven dataset.
Hot songs: We select two non-stop playlists of hot songs, featuring top-ranking Billboard songs from 2019 and 2022(YoutubeMix, 2019, 2022), spanning various genres like Pop, Rock, and Hip-hop. Both playlists are around two hours long, sourced from YoutubeMix adhering to the same format.
Bass: To ensure the inclusion of low-frequency audio, we had two bass playlists (YoutubeMix, 2024, 2023a), lasting for one hour and two hours, both of which are downloaded from YoutubeMix under the same data format as above.
Speech: Besides musical audio, we add two conversational speech datasets. They are a mental health podcast featuring Selena Gomez(YoutubeMix, 2023c) and a Conan O’Brien podcast(YoutubeMix, 2023b). Both last over one hour long and are downloaded by YoutubeMix in the same format.
All datasets are then cut into equal-length clips, paired with same-length sensing signal clips as defined in section 3.3.
4.2. Experiment results
@@
Figure 5. The model generalizes well to unseen datasets and new types of music or speech.

### v3

@@ III-A Understanding the effect of dynamically scaling the signal
Achieving our two objectives simultaneously is hard because tweaking signal amplitude in time domain will distort frequency domain. To understand the effect of inordinately tweaking the amplitude, consider amplitude modulation (AM) as an interpretable example, a telecommunications technique to transmit information by varying carrier wave amplitude. In Fig. 2, the high-frequency carrier is multiplied by the low-frequency base signal, i.e. amplitude modulated, making the frequency spectrum spread out with noisy side bands. In this case, the change is closed-form; side frequencies are $f_{c}+f_{m}$ and $f_{c}-f_{m}$ (Appendix A). Building upon this observation, our cognitive scaling is namely multiplying the signal with a more complex window function, leading to various sidebands in the frequency. Therefore, this effect of altering signals necessitates our utilization of deep learning; closed-form signal processing or traditional machine learning can hardly dynamically control the distortion during generation.
III-B Sensing signal modulation types
@@
Acoustic sensing commonly uses two types of signals: sine waves and FMCW chirps, which serve different purposes in various applications. 1) Sine waves are single-frequency oscillation that varies with time according to the mathematical function $A\sin(2\pi ft+\phi)$, where $A$ is the amplitude of the sine wave, $f$ is the frequency, $\phi$ is the initial phase. They are common for tasks such as signal testing, calibration, and reference signals, providing a well-defined waveform, easy to generate and analyze. 2) FMCW chirps have frequency changes linearly over time, basically repeated chirps defined as $A\sin(2\pi(f_{a}t+\frac{f_{1}-f_{0}}{2T}t^{2}))$, where $f_{0}$ and $f_{1}$ are the starting and ending frequency; $T$ is the duration per chirp. They are widely used in radar and sonar, offering improved range resolution for precise distance measurements and multi-target discrimination.
III-C Define the problem as optimization
@@
Input and output: Let $x$ be the sensing signal, i.e. a high-frequency sine or chirp. Concurrently, an App plays music $z$. Then the mixer mixes them within the maximum range of [m,m][-m,m] allowed by the number of DAC bits. In this paper, we use signed 16-bit PCM, resulting in a range of $2^{15}\sim 2^{15}-1$. For computation convenience, we set $m=2^{15}-1$. In general, our goal is to generate a cognitively-scaled sensing signal $\hat{x}$ conditioned on the music $z$; and the $\hat{x}$ is optimal when satisfying the following objectives defined in the loss.
Loss: The first objective is to maximize the sensing amplitude within the bandwidth left by music. In other words, the amplitude of the mixed signal $\hat{x}+z$ in time domain should be close to the maximum allowed amplitude $m$. To restate, it is to minimize loss $q(\hat{x}+z,m)$, namely the difference between the magnitude of them. However, this would cause consequential frequency distortion as explained in section III-A, which deteriorates the receiver-side algorithms since they generally rely on accurate frequency (either FFT convolution or correlation convolution). Therefore, we define another loss p(x,x^)p($\hat{x}$) as the difference between $x$ and $\hat{x}$ in frequency domain; it is the difference between their L-2-norm of FFT coefficients. In summary, our goal is to minimize these two loss terms simultaneously as an optimization problem.
@@
$p(x,\hat{x})=1-\|\hat{c_{f_{c}}}\|_{2}+\frac{1}{N-1}\sum_{i\in-{f_{c}}}\|c_{i}-\hat{c_{i}}\|_{2}$ (1)
[…]
$q(\hat{x}+z,m)=1-\frac{1}{2N}\|\frac{\hat{x}+z}{m}\|_{2}$ (2)
For FMCW chirp, the multiple frequency bins also need even amplitude to correlate well. So, an additional variance-loss for FMCW is built via standard deviation with zero correction:
$s(\hat{x})=\sqrt{\frac{\sum_{i\in{f_{c}}}(\hat{c_{i}}-\bar{\hat{c_{i}}})^{2}}{N_{f_{c}}-1}}$ (3)
@@
$loss_{sine}$ $=\alpha*p(x,\hat{x})+\beta*q(\hat{x}+z,m)$ (4)
[…]
We will ablate study the hyperparameters in section IV. Note that, our loss may not guarantee that $\hat{x}+z$ falls within [m,m-m,m]. Therefore, we need to embed this constraint into model layers, as introduced in the following section III-D per customized link function. In practice, we split the sensing signal and music into small windows and run optimization for each window. For chirps, the window size should match the chirp size. In this paper, we use 18KHz sine wave and 18k-20kHz FMCW signal, with a window size of 512 and a sampling rate of 48KHz. These are common settings balancing efficiency and performance. And our model can be extended to other settings as long as training and testing settings match.
III-D Cognitive scaling neural network
@@
Model architecture with WaveNet backbone. To elaborate, it combines a non-auto-regressive sequence-to-sequence structure with WaveNet backbones. While WaveNet architecture was originally developed for AR modeling, it has since become a basic backbone of various audio models. And we validate this choice in the micro-benchmark section by experimental comparison with sample CNN and RNN backbones. In detail, it employs a stack of dilated convolutional layers which provides better reception fields than conventional ones, making it adept at raw waveform modeling and processing sequential data more efficiently than RNNs. The input of the model is concatenated sensing signal $x$ and music $z$ of the same length of window size 512. Then the 2-row tensor is passed through 2 layers of Wavenet with fused addition and tanhtanh + multiplication, then dilated convolution (hidden size of 2, kernel size of 5, and dilation rate of 1), and finally residual skip layer (double hidden size except the last layer). Then the outputs pass through conv1d layer to merge the two rows, followed by custom layers to be described in the following sub-sections. Based on observation, we choose a large learning rate of 0.1 with Adam optimizer that is stable in training. Our implementation refers to DiffSinger [30] on HuggingFace.
Figure 3: Cognitive scaling neural network.
@@
Figure 4: The model generalizes well to unseen datasets and new types of music or speech.
A sinc kernel for anti-aliasing. In early experiments, we observe non-negligible low-frequency noise in the model output, especially when modeling FMCW chirps. It is a common problem in computer vision community that the neural network learns low frequency faster and also down-scaling in layers like convolution and max pooling could induce aliasing. Therefore, we design a custom sinc kernel for band-passing middle activations to isolate specific frequency bands of interest within the signal. The unique properties of the (sinc) filter, notably its brick-wall characteristic in the frequency domain as its Fourier transform is a rectangular function (Appendix  B), makes it a preferred choice of filters over Hamming, etc. So, by putting (sinc) kernel as a layer before the final link function, it mitigates artifacts that may arise in the first stage. For implementation, we use sinc_impulse_response function in the Pytorch experimental nightly version. The negative of it is a high-pass filter. And they can sum to one band-pass filter or more. To demonstrate its effectiveness, we perform ablation in the micro-benchmark section. Another advantage is its potential to serve as a parameterized kernel when building models for new types of sensing signals other than chirp or sine. A parameterized cutoff could learn a dynamic filter for even multiple signals trained in one model. While we currently train standalone models dedicated to each sensing signal, so that each model achieves optimal performance since we focus on substantiating and validating the underlying concept.
A link function for normalization with agnostic-music as bias. As described, $\hat{x}+z$ should be within the range of [m,m-m,m], namely a normalization with agnostic music as bias. To restate, the output $\hat{x}$ should be $a\times m+z$ where $a$ is the activation from the previous layer within [0, 1]. So, to embed this constraint as a trainable layer, we leverage the concept of link function in Generalized Linear Models (GLM); It is in analogy to the logit function for the logistic regression, which uses the prior to target on the probability of binary classification ranging from 0 to 1. Similarly, CoPlay uses a hyperbolic tangent function (tanhtanh) as link function that scales the output within [-1, 1]. A bonus is that it encourages the value to approach -1 or 1, which is useful in our cases because we want to maximize the magnitude of the waveform, i.e. close to 1 or -1. In contrast, clipping simply truncates the output, resulting in loss of information or under-fitting, while tanhtanh is more effective especially when data is continuous and has a wide range of values. So, we design a tanhtanh layer rather than a min-max clipping. $\hat{x}=tanh(a)*m-z=\frac{e^{a}-e^{-a}}{e^{a}+e^{-a}}*m-z$ where $a$ represents the activations from the previous layer. This scales $\hat{x}+z$ within [m,m][-m,m]. For inference, we feed $\hat{x}+z$ into the mixer in case $\hat{x}$ exceeds the numerical range. Finally, to quantify the output as integers while the model is trained in float64, we employ a workaround using the round function: it detaches the difference and then adds it back. Otherwise, round or typecast are not differentiable in PyTorch.
IV Micro-benchmark
This section details the experiment setup and various model performance micro-benchmarks, including different datasets, music types and volume. Ablation studies also justify our design in model architecture and multi-task learning weights. Code is published on Github [^1] https://github.com/lydhr/CoPlay.
Evaluation metric in this section is mainly the loss, as defined in section III-C. It indicates the relative performance among micro-benchmarks. Furthermore, to offer interpretable analysis of output quality, we provide visualizations for frequency spectrograms and cross-correlation ranging profiles. In the next section, we will further assess the performance in downstream field studies like gesture and respiration detection.
IV-A Datasets
We use four audio types, each with two datasets, to evaluate generalization across agnostic audio. They are cut into equal-length clips, then paired with same-length sensing clips as defined above. 1) Piano: One dataset is the Beethoven dataset [25], consisting of 10-hour piano sonatas, often serving as a benchmark music dataset for audio generation. We interpolate the data to a sample rate of 48kHz and cast the 8-bit quantization into a 16-bit signed integer. Another dataset is the YouTubeMix dataset [31] with 4-hour higher-quality recordings, which we format in the same way as the Beethoven dataset. The following are also downloaded via YoutubeMix toolset adhering to the same format. 2) Hot songs: We select two non-stop playlists of hot songs, featuring top-ranking Billboard songs from 2019 and 2022[32, 33], spanning various genres like Pop, Rock, and Hip-hop. Both are around two hours long. 3) Bass: To ensure the inclusion of low-frequency audio, we had two bass-centric playlists [34, 35], lasting for one hour and two hours. 4) Speech: Besides musical audio, we add two conversational speech datasets: a podcast featuring Selena Gomez[36] and a Conan O’Brien podcast[37]. Both are over one hour.
IV-B Experiment results
Generalization across datasets and music types: In total, we have 8 datasets. We define intra-dataset experiment as training and testing the model with the same dataset, cross-dataset as training and testing with same-type different datasets, and cross-type as with different types of datasets (the most out-of-distribution case). Note that, we avoid randomly splitting the dataset for testing that is non-realistic in usage. Instead, the last 10% sequentially is testing set. The rest is then randomly split into 80% training and 20% validation. All trainings run 10 epochs to converge unless otherwise specified. Fig. 4 shows the average error for sine and chirps separately, grouped by music types. For better plots, the y-axis is log-scaled testing loss multiplied by $1e5$. In summary, the results demonstrate strong generalization, as the errors in cross-dataset and cross-type are similar. Intra-dataset is the best as expected because the distribution is close between testing and training. However, for chirp data, intra-dataset is counter-intuitive, and it takes 10x more epochs to converge. This probably owes to the complexity of chirp that needs more training data to learn effectively, while intra-dataset uses only 6̃0% of the data available in cross-dataset/type setup. Notably, speech also achieves robust cross-type results, highlighting great generalization ability, as the most distinguishable type. Overall, errors remain low, which are further justified by downstream tasks in following section.
target
-loss
recovery
-loss
amplitude
-loss
variance
-loss
sine 0.5928 0.0012 0.9558 -
chirp 0.1852 0.0013 0.9560 0.0214
TABLE I: The value of sub-loss terms.
Decompose the loss: The overall error comprises multiple loss terms as defined in section III-C. Hence, we decompose it into individual loss term values to assess each corresponding objective. Table. I is the average sub-loss on hot songs datasets. They are training loss including three sub-loss values for sine and an additional variance-loss for FMCW as defined above. By default weights, each sub-loss ranges from 0 to 1. We observe that target-loss and amplitude-loss are the bottlenecks, aligning with the intuition that minimizing frequency distortion and maximizing the amplitude are opposed objectives, which is a common issue in multi-task learning. This table serves as a reference for tweaking the weights of loss terms to specific downstream task requirements. For instance, near-field sensing tasks can sacrifice amplitude for minimal distortion.
Varying music volume: Although a high music volume might degrade our model performance, Fig. 5(b) shows that the overall error does not increase with music volume. In detail, we train with the same sub-loss experiment and test with increasing levels of music volume. The x-axis indicates the ratio of sensing signal volume and music volume. The default ratio is 1:1, and 0.6:1.4 is when music is roughly double in power, equivalently 7.3db louder. So, CoPlay is stable when training with agnostic music even in unseen volume, likely due to the normalization in amplitude-loss and recovery-loss.
Ablation of model architecture: To justify our choice, we compare three models backbones namely CNN, RNN, and Wavenet with the same number of recursive layers. Fig. 5(a) shows that Wavenet slightly outperforms CNN and both surpass RNN. But CNN has a limited reception field so we chose Wavenet, which improves this with dilated convolution layers. Furthermore, the sample RNN uses LSTM layers, making it less efficient to train.
Ablation of multi-task loss terms: Our loss function comprises sub-losses with equal default weights, which can be adjusted to emphasize specific objectives (as in  III-C). So, we test increasing recovery-loss weight from 1 to 1.2, 1.4, and 1.6 to prioritize minimizing frequency distortion, while the other weights are 1 by default; this is a practical use case when certain sensing systems, like near-field tasks, can compromise the power for minimal frequency distortion. Fig. 5(c) shows error decreasing as weight increases, providing a reference for tuning this ratio for the downstream tasks in practice.
Ablation of sinc filter: To validate the efficacy of the sinc filter, Fig. 5(d) exhibits improved performance with the presence of sinc kernel, in both the sine and chirp models, especially the former. We attribute this enhancement to the narrow target main lobe in the spectrum of sine, thus more side lobes are to be suppressed. The sinc kernel helps significantly reduce these lobes, resulting in a lower numerical loss. So, we incorporate this essential layer in all experiments.
IV-C Spectral analysis
Training procedure in 3D frequency domain: Although the numerical loss can indicate the model performance, we seek a more comprehensive perspective by visualizing the model output in frequency spectrum. This helps confirm that the output aligns well with expectations. In Fig. 6, we take the FFT of a sample window output $\hat{x}$ and screenshot it with the increasing training epochs. The rise of the yellow peaks indicates that it converges to an 18kHz sine wave or 18k-20kHz FMCW as the training epoch increases. The side lobes are much weaker than the main lobes, which manifest good SNR. (Moreover, we further substantiate the SNR by qualitative study in section V)